OpenStudy (anonymous):

A company's sales are seasonal with the peak in mid-December and the lowest point in mid-June. The company makes $100,000 in sales in December, and only $20,000 in June. (a) Find a trigonometric function, s = f(t), representing sales at time t months after mid-January i know the period is 12 and the midline is 60,000 and the amp is 40,000 but i don't know where to go from there

OpenStudy (anonymous):

\[60000+40000\cos(\pi/6t)\] that is what i have so far but I know that is wrong

OpenStudy (baru):

\(60000+40000cos (\frac{\pi }{6}t -\phi) \)

OpenStudy (baru):

find \(\phi\) by plugging in t=6, f(t)=20000